I have to show the five steps...
-> Two brothers are saving money to buy tickets to a concert. Their combined savings is $55. One brother has $15 more than the other. How much has each saved?

Answer by stanbon(75887)

Two brothers are saving money to buy tickets to a concert. Their combined savings is $55. One brother has $15 more than the other. How much has each saved?
-------------------------------
Let the two brothers savings be a and b.
a + b = 55
a = b + 15
-------
Substitute for "a" to solve for "b":
b+15 + b = 55
2b = 40
b = 20 (amt saved by b)
------------------------------
But a + b = 55, so a = 35
==================================
Cheers,
Stan H.
Reply to stanbon@comcast.net